OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3.

OpenStudy (anonymous):

will give medal to best answer

OpenStudy (texaschic101):

y = mx + b m = 3 (4,2) x = 4 and y = 2 sub in your stuff 2 = 3(4) + b 2 = 12 + b 2 - 12 = b -10 = b y = 3x - 10

OpenStudy (haseeb96):

y2-y1=m(x2-x1) m=3 (x2,y2)=(4,2) 2-y1=3(4-x1) 2-y1=12 -3x1 3x1-y1+2-12=0 3x -y -10 =0

OpenStudy (haseeb96):

this is the correct equation

OpenStudy (texaschic101):

sorry about what I posted and then deleted before...I wrote point slope formula instead of slope intercept

OpenStudy (texaschic101):

slope intercept form is y = mx + b your correct equation is : y = 3x - 10